[figure: 9.7 cm]

THE LOSS OF MAGNETISM IN FeS UNDER HIGH PRESSURES

Z. Ropka^{a} and R.J. Radwanski^{a,b} aCenter of Solid State Physics, Snt Filip 5, 31-150 Krakow,Poland. bInstitute of Physics, Pedagogical University, 30-084 Krakow, Poland

We argue that the formation of the non-magnetic state of FeS in external pressures above 6.5 GPa is in agreement with the prediction of the quantum atomistic solid-state theory (QUASST), an extension of the crystal-field theory with the spin-orbit coupling incorporated, and is related to the change of sign of the off-cubic trigonal distortion marked by the change of the c/a ratio from > 1.63 (magnetic) to < 1.63 (non-magnetic). In the calculations we take into account very strong correlations among six d electrons within the incomplete 3d shell, the crystal field, relativistic spin-orbit coupling and two Hund’s rules. Our QUASST approach allows to calculate both zero-temperature properties and whole thermodynamics.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : R.Radwanski

Address for correspondence : Center of Solid State Physics, Snt Filip 5, 31-150 Krakow, Poland

Email address : sfradwan@cyf-kr.edu.pl